Sevall Inc's Mexican Subsidiary, Cancunrocks, is expected to pay Sevall 1 Peso per share in 1 year after all foreign and US taxes have been accounted for. The exchange rate in 1 year is expected to be $0.10 dollars per peso. After this, the peso is expected to depreciate against the dollar by 2% forever due to inflation differences between the US and Mexico. The peso denominated dividend is expected to grow at 7% a year indefinitely. Sevall owns 10M shares of Cancunrocks. Assuming that the cost of equity is 12%, what is the valuation of Cancunrocks in US dollars?
Answer #1

=10*(1*0.1/1.12+1*1.07*0.10*(1-2%)/1.12^2....)
=10*(1*0.1/1.12*(1+1.07*(1-2%)/1.12.....))
=10*(1*0.1/1.12*1/(1-1.07*0.98/1.12))
=14.00560224 million US dollars